Solve for x & y
x/2 + y/2 = 9 and
sqrt of xy = 7.2

Answer by mananth(16946)

x/2 + y/2 = 9 and
x+y=18
sqrt of xy = 7.2=7 1/5 = 36/5
xy = 1296/25
y = 1296/25x
plug value of y in the other equation
x+1296/25x=18
25x^2+1296=450x
25x^2-450x+1296=0
Solve using quadratic formula
Discriminant =72900

x=14.4 OR 3.6
y=3.6 OR 14.4